I'm trying to determine a quadratic equation that exhibits such:
The equation I am working with is the expected number of ticket sales for each day where x=1.
Tickets = -0.2x2 + 12x + 11
I need to find a quadratic equation to determine the last day that tickets will be sold. (Note: Write your answer in terms of the number of days after ticket sales begin.)
I have tried: -0.2x2 + 12x + 11 = t (representing number of tickets)
Should I find t first and then substitute t to find x (number of days)? Or should I have the equation equal 0? (-0.2x2 + 12x + 11) = 0
I would really appreciate your help with this problem.
Thank you

Answer by stanbon(75887)

I'm trying to determine a quadratic equation that exhibits such:
The equation I am working with is the expected number of ticket sales for each day where x=1.
Tickets = -0.2x2 + 12x + 11
I need to find a quadratic equation to determine the last day that tickets will be sold. (Note: Write your answer in terms of the number of days after ticket sales begin.)
I have tried: -0.2x^2 + 12x + 11 = t (representing number of tickets)
Should I find t first and then substitute t to find x (number of days)? Or should I have the equation equal 0? (-0.2x^2 + 12x + 11) = 0
-----------------------------------------------
I think you want to solve -0.2x^2 + 12x + 11 > 0
That will tell you for what values the number of tickes sold is positive.
------------------------
So solve -0.2x^2 + 12x + 11 = 0
x = [-12 +- sqrt(144 - 4*-0.2*11)]/(2*-0.2)
x = [-12 +- sqrt(152.8)]/(-0.4)
x = [-12+-12.36]/(-.4)
x = 0.36/-0.4 = -0.9
OR
x = (-12-12.36)/(-.4)
x = 60.9
-------------------
So the quadratic is positive for -0.9 < x < 60.9
The last day tickets are sold is day # 60
The 1st day tickets are sold is day # 0
==============================================
Cheers,
Stan H.